OpenStudy (anonymous):

What is the distance between the points (4, 7) and (4, –5)? A. 0 units B. 4 units C. 2 units D. 12 units

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\) does this help ??

OpenStudy (anonymous):

No

OpenStudy (anonymous):

I beleive the answer is D 12 Units. 4-4=0^2=0 and 7+5=12^2=144 Squared=12 Units D

OpenStudy (anonymous):

Yea me too

OpenStudy (anonymous):

What is the distance between the points (1, –6) and (–5, 2)? Round your answer to the nearest tenth. A. 7.2 units B. 10.0 units C. 5.6 units D. 9.0 units

OpenStudy (anonymous):

Can someone help me?

OpenStudy (anonymous):

Does anyone have skype so they can help me?

OpenStudy (anonymous):

1+5=6^2=36 + -6 -2=-8^2=64 36+64=128 Squared =10 Units. Mind hitting that button to help me out on my points thanks

OpenStudy (anonymous):

Which button?

OpenStudy (anonymous):

Best response

OpenStudy (anonymous):

Who has skype?